### v1

# Safe Flow Matching: Robot Motion Planning with Control Barrier Functions

Abstract
Recent advances in generative modeling have led to promising results in robot motion planning, particularly through diffusion and flow-based models that capture complex, multimodal trajectory distributions. However, these methods are typically trained offline and remain limited when faced with unseen environments or dynamic constraints, often lacking explicit mechanisms to ensure safety during deployment. In this work, we propose, Safe Flow Matching (SafeFM), a motion planning approach for trajectory generation that integrates flow matching with safety guarantees. By incorporating the proposed flow matching barrier functions, SafeFM ensures that generated trajectories remain within safe regions throughout the planning horizon, even in the presence of previously unseen obstacles or state-action constraints. Unlike diffusion-based approaches, our method allows for direct, efficient sampling of constraint-satisfying trajectories, making it well-suited for real-time motion planning. We evaluate SafeFM on a diverse set of tasks, including planar robot navigation and 7-DoF manipulation, demonstrating superior safety, generalization, and planning performance compared to state-of-the-art generative planners. Comprehensive resources are available on the project website: https://safeflowmatching.github.io/SafeFM/
Keywords: Flow Matching, Safe Motion Planning, Control Barrier Functions, Trajectory Generation
1 Introduction
Enabling robots to autonomously generate safe and feasible motions in dynamic environments remains one of the central challenges in robotic learning and motion planning. Recent advances in deep generative modeling, particularly in diffusion models, have significantly expanded the capabilities of robots to learn diverse and high-dimensional behaviors from demonstrations or interactions. These models have proven effective in modeling multimodal trajectory distributions, enabling robots to adaptively generate context-dependent motions across various tasks, such as navigation [1], object manipulation [2], and locomotion [3].
Despite their success, diffusion models come with notable limitations. The inference process is typically slow and computationally intensive, as it involves iteratively solving a stochastic differential equation (SDE) over many denoising steps. This high inference cost impedes real-time decision making, especially in safety-critical applications where rapid reactions to environmental changes are necessary. Furthermore, while diffusion models can be conditioned on goals or task descriptions, they generally lack formal safety guarantees, which is crucial when deploying robots in dynamic environments with obstacles, humans, or other agents. Recent works have begun to address these concerns by incorporating safety constraints into diffusion-based frameworks (e.g., via model-predictive control or barrier functions), but doing so often further increases computational complexity and may compromise sample quality or task performance.
To overcome these challenges, we turn to Flow Matching (FM)—a recently emerging generative modeling framework that offers a promising alternative to diffusion models. Instead of relying on stochastic processes, FM defines a deterministic flow field—governed by an ordinary differential equation (ODE)—that smoothly transforms samples from a simple prior distribution into target data samples. FM has demonstrated strong performance across several domains, including image synthesis, video generation, speech, proteins, and robot trajectory learning, largely due to its numerical stability, training simplicity, and efficient inference. More importantly, the FM framework is amenable to explicit constraint integration, making it a natural candidate for generating safe robot motions.
In this work, we introduce Safe Flow Matching (SafeFM), a novel method for constraint-aware motion planning that combines the efficiency and flexibility of flow matching with the formal safety guarantees provided by Control Barrier Functions (CBFs). SafeFM learns a continuous-time vector field that maps random waypoints from a prior distribution to goal-directed robot trajectories, while simultaneously ensuring that the generated trajectories respect predefined safety constraints throughout the entire flow. By leveraging finite-time flow invariance principles, we embed safety specifications directly into the learned dynamics, enabling real-time generation of safe and feasible motions without sacrificing generation quality or expressiveness.
To evaluate the efficacy of SafeFM, we conduct extensive experiments across several simulated and real-world scenarios, including planar robot navigation, 7-DoF robotic arm manipulation, and environments with unseen obstacle configurations. We compare SafeFM to both diffusion-based approaches and unconstrained flow-based baselines, demonstrating that our method consistently achieves higher success rates, lower violation of safety constraints, and significantly faster inference times. The main contributions of this paper are:
  • We propose SafeFM, a novel flow matching framework for robot motion planning under safety constraints.
  • We introduce a theoretically grounded integration of control barrier functions into flow-based generative models to ensure safe trajectory generation.
  • We empirically demonstrate the superior safety, efficiency, and generalization capabilities of SafeFM across multiple tasks.
2 Related Work
Generative models for robot learning have gained considerable attention in recent years. This section reviews recent advancements in diffusion-based and flow-based methods, with a focus on their applications to robot learning.
Diffusion-based Models for Robot Learning.
Diffusion models have been effectively utilized in various robot learning domains, including motion planning, imitation learning, and reinforcement learning, as demonstrated by works such as [4, 2, 5]. In imitation learning, diffusion models have been leveraged to produce robot trajectories based on expert demonstrations. For example, Pearce et al. [6] and Reuss et al. [5] employ diffusion models to develop policies that replicate expert behaviors. Additionally, Xian et al. [7] introduce ChainedDiffuser, a novel policy architecture that integrates action keypose prediction with trajectory diffusion generation for robotic manipulation tasks. To tackle long-horizon planning and task execution, Li et al. [8] and Mishra et al. [9] propose skill-centric diffusion models that combine learned distributions to create extended plans during inference. For generating control actions, Chi et al. [10] present Diffusion Policy, which frames policy learning as a conditional denoising diffusion process over the robot’s action space, conditioned on 2D observation features. This concept is advanced by Ze et al. [11] in their 3D Diffusion Policy, which incorporates complex 3D visual representations to enable generalizable visuomotor policies from just a few dozen expert demonstrations.
Recent efforts have also aimed to ensure that generated trajectories comply with the physical laws governing robot dynamics. For instance, Gadginmath and Pasqualetti [3] embed system dynamics directly into the denoising process, while Chen et al. [12] use generated state sequences with an inverse dynamics model to derive actions for coordinated bimanual tasks. To enforce safety and dynamical constraints, Li et al. [13] incorporate penalties for constraint violations into the loss function, though this method does not guarantee strict adherence to hard constraints. Alternatively, Bouvier et al. [14] project generated policies onto a dynamically admissible manifold during training and inference. However, this technique underapproximates the reachable set using a polytope, relying on the assumption of smooth nonlinear dynamics and environments. Drawing inspiration from control barrier functions, Xiao et al. [15] propose Safe Diffuser to enforce safety constraints within diffusion models, influencing our focus on safety in robot motion generation. Furthermore, Mizuta and Leung [16] integrate both system dynamics and control barrier functions into the diffusion framework.
Despite these advances, diffusion models struggle in high-dimensional control tasks, such as robotic manipulation, where achieving real-time inference with high-quality actions remains difficult. This challenge stems from the inherent uncertainty and iterative nature of the diffusion process.
Flow-based Models for Robot Learning.
Flow matching models have emerged as a promising alternative to diffusion models, offering potential improvements in efficiency and performance for robotic applications. For instance, Zhang and Gienger [17] introduce a flow matching policy that utilizes affordance-based reasoning for manipulation tasks. Departing from image-based inputs, Chisari et al. [18] employ point clouds with conditional flow matching to generate robot actions. To enhance efficiency, Zhang et al. [19] propose a flow policy that uses consistency flow matching with a straight-line flow, enabling one-step generation and striking a balance between speed and effectiveness, though limitations persist in complex manipulation scenarios. Capitalizing on the strengths of flow matching, Braun et al. [20] apply these models to robot states defined on a Riemannian manifold. This approach better captures high-dimensional multimodal distributions, improving the handling of complex robotic systems.
3 Preliminaries
3.1 Flow Matching Model
A flow matching model is defined through an ODE governed by a time-dependent vector field $\bm{v}^{\bm{\theta}}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ parameterized by a neural network, which induces a time-dependent flow map $\bm{\psi}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, i.e.,
$\frac{\mathrm{d}}{\mathrm{d}t}\bm{\psi}_{t}(\bm{\tau})=\bm{v}_{t}^{\bm{\theta}}(\bm{\psi}_{t}(\bm{\tau})),$ (1)
where $\bm{\psi}_{t}(\bm{\tau}):=\bm{\psi}(t,\bm{\tau})$ and $\bm{v}_{t}^{\bm{\theta}}:=\bm{v}(t,\bm{\tau})$ with initial conditions $\bm{\psi}_{0}(\bm{\tau}_{0})=\bm{\tau}_{0}$. We represent a trajectory $\mathcal{T}(t):[0,1]\to\mathbb{R}^{H\times d_{s}}$ as a sequence of states $\mathbf{s}_{t}^{k}\in\mathbb{R}^{d_{s}}$, such that $\mathcal{T}_{t}=[(\mathbf{s}_{t}^{0})^{\top},\dots,(\mathbf{s}_{t}^{H})^{\top}]\in\mathbb{S}^{H+1}\subseteq\mathbb{R}^{(H+1)d_{s}}$, with the dimensionality $d=d_{s}(H+1)$. Considering a start $\mathcal{T}_{0}=\bm{\tau}_{0}$ at $t=0$, the vector field generates the probability path $p_{t}$ if its flow $\bm{\psi}_{t}(\bm{\tau})$ satisfies
$\mathcal{T}_{t}:=\bm{\psi}_{t}(\mathcal{T}_{0})\sim p_{t}~{}\text{for}~{}\mathcal{T}_{0}\sim p_{0}$ (2)
Thus, the velocity field $\bm{v}_{t}^{\bm{\theta}}$ enables sampling from the probability path $p_{t}$ by solving the ODE. The objective of flow matching is to train a parameterized vector field $\bm{v}t^{\bm{\theta}}$ such that its induced flow $\bm{\psi}t$ generates a probability path $p_{t}$, transitioning from an initial distribution $p_{0}=p$ to the target data distribution $p_{1}=q$.
3.2 Safe Invariance in Dynamical Systems
Consider a control-affine system
$\dot{\bm{x}}_{t}=\bm{f}(\bm{x}_{t})+\bm{g}(\bm{x}_{t})\bm{u}_{t},$ (3)
where $\bm{x}\in\mathbb{X}\subset\mathbb{R}^{n}$, $\bm{f}(\cdot):\mathbb{X}\to\mathbb{R}^{n}$ and $\bm{g}(\cdot):\mathbb{X}\to\mathbb{R}^{n\times m}$ are locally Lipschitz, and $\bm{u}\in\mathbb{R}^{m}$ is the control input. Moreover, a super level set :=$\mathbb{C}:=\{\bm{x}_{t}\in\mathbb{R}^{n}|h(\bm{x}_{t})\geq 0\}$ is defined via a continuously differentiable equation $h(\cdot):\mathbb{R}^{n}\to\mathbb{R}$, which encode the safety guarantees for Eq. 3.
Definition 1 ([21]).
Given a safe set $\mathbb{C}$, $h(\bm{x})$ is a time-invariant CBF for system Eq. 3 if there exists a class $\mathcal{K}$ function $\alpha(\cdot)$ such that
$\sup_{\bm{u}\in\mathbb{R}^{m}}\{L_{\bm{f}}h(\bm{x})+L_{\bm{g}}h(\bm{x})\bm{u}+\alpha(h(\bm{x}))\}\geq 0$ (4a)
for all $\bm{x}\in\mathbb{C}$, where $L_{\bm{f}}h(\bm{x}),L_{\bm{g}}h(\bm{x})$ denote the Lie derivatives of $h(\bm{x})$ along $\bm{f}(\bm{x})$ and $\bm{g}(\bm{x})$, respectively.
The safety guaranteed by the control barrier function is shown as follows.
Theorem 1 ([21]).
If $h(\cdot)$ is a valid control barrier function and $\bm{x}(0)\in\mathbb{C}$, then $\mathbb{C}$ is forward invariant indicating $\bm{x}(t)\in\mathbb{C}$ for $\forall t\in\mathbb{R}_{0,+}$.
However, the condition of initial safety $\bm{x}_{0}\in\mathbb{C}$ may not be satisfied in some scenarios. Particularly, in robotic motion planning tasks, it is desired that the generated trajectory enters the safe set $\mathbb{C}$ within a prescribed time $T_{d}\in\mathbb{R}_{+}$, and remains in $\mathbb{C}$ afterwards. Thankfully, this task can be defined as prescribed-time safety, i.e.,
Definition 2.
A system in Eq. 3 is prescribed-time safe with a predefined time $T_{d}\in\mathbb{R}_{+}$, if $\bm{x}_{t}\in\mathbb{C}$ for any $t\geq T_{d}$ with arbitrary $\bm{x}_{0}\in\mathbb{X}$.
To ensure safety in flow matching, we draw inspiration from the safety guarantees provided by control barrier functions and propose an adaptation for trajectory generation. The mathematical formulation of this approach is presented in the following section.
4 Safe Flow Matching
The objective of safe flow matching is to generate a trajectory $\mathcal{T}=[(\mathbf{s}^{0})^{\top},\dots,(\mathbf{s}^{H})^{\top}]^{\top}$ that closely aligns with a target distribution while remaining within a predefined safe set. In this section, we formally define safe flow matching and introduce the concept of flow barrier functions. Furthermore, we propose safety regularization terms designed to enforce compliance with safety constraints during trajectory generation.
4.1 Safe Flow Matching under Unified Constraints
In this subsection, we consider a safe set $\mathbb{C}_{s}\subseteq\mathbb{S}$ defined for each state $\mathbf{s}^{i}$, $i=0,\dots,H$, characterized by a single continuous function $h(\cdot):\mathbb{S}\to\mathbb{R}$ such that The generated trajectory $\mathcal{T}$ is called as safe, if all generated states $\bm{s}^{i}$ remain in the safe set $\mathbb{C}_{s}$, which is mathematically defined as follows
Definition 3 (Trajectory Safety).
A trajectory $\mathcal{T}$ with length $H$ is safe w.r.t $h(\cdot)$, if
$h(\bm{E}_{i}\mathcal{T})\geq 0,$ $\forall i\in\{0,1,\cdots,H\}$ (5)
where $\bm{E}_{i}\in\mathbb{R}^{d_{s}\times d}$ is a selection matrix defined as $\bm{E}_{i}=[\bm{0}_{d_{s}\times id_{s}},\bm{I}_{d_{s}},\bm{0}_{d_{s}\times(H-i)d_{s}}]$.
As the trajectory $\mathcal{T}$ is generated via a flow model in finite time, the trajectory safety can be guaranteed. Moreover, unlike conventional CBFs, flow matching must account for potentially unsafe initial trajectories $\mathcal{T}_{0}\sim p_{0}$. Such a safe trajectory generation task in flow matching is defined as a prescribed-time safety problem, where the initial state is required to start in the predefined safe set. Accordingly, we introduce the definition of safe flow-based trajectory generation as follows.
Definition 4 (Safe Flow Matching).
A flow matching process is called safe if the generated trajectory $\mathcal{T}_{1}$ at $t=1$ satisfies the safety conditions in Definition 3 for any initial sample $\mathcal{T}_{0}\sim p_{0}$.
To ensure the safety of the generated trajectory as in Definition 4, we consider the control input $\bm{u}_{t}=[{\bm{u}_{t}^{0}}^{\top},\cdots,{\bm{u}_{t}^{H}}^{\top}]^{\top}$ serving as a regularization term to steer the flow toward a safe manifold. Therefore, the flow model is formulated as
$\frac{\mathrm{d}}{\mathrm{d}t}\bm{\psi}_{t}(\bm{\tau})=\bm{v}_{t}^{\theta}(\bm{\tau})+\bm{u}_{t},~{}\forall t\in[0,1].$ (6)
Drawing inspiration from control barrier functions, we propose a flow matching barrier function (FMBF) to enforce safety constraints in the flow matching framework, which is defined as
Definition 5 (Flow Matching Barrier Function).
A function $h(\cdot)$ is called a flow matching barrier function if it satisfies the following conditions:
  1. 1.
    For all $i=0,\cdots,H$ and $t\in[0,1)$, it has
    $\sup_{\bm{u}_{t}^{i}\in\mathbb{R}^{d_{s}}}\Big{\{}\frac{\mathrm{d}h(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\mathrm{d}h(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{u}_{t}^{i}+\varphi(t,h(\bm{E}_{i}\mathcal{T}_{t}))h(\bm{E}_{i}\mathcal{T}_{t})\Big{\}}\geq 0,$ (7)
    where
    $\varphi(t,h(\bm{s}))=\begin{cases}\varphi_{0},&\text{if~{}}h(\bm{s})\geq 0\\ \varphi_{1}(t),&\text{otherwise}\end{cases}$ (8)
    with $\varphi_{0}\in\mathbb{R}_{+}$ and a blow-up function $\varphi_{1}(\cdot):[0,1)\to\mathbb{R}_{+}$ satisfying $\lim_{t\to 1^{-}}\varphi_{1}(t)=+\infty$.
  2. 2.
    For all $i=0,\cdots,H$, the terminal condition holds: $h(\bm{E}_{i}\mathcal{T}_{1})\geq 0$.
Build upon the FMBF, the safety guarantee derived in the following theorem.
Theorem 2.
Suppose $h(\cdot)$ is a valid flow matching barrier function as per Definition 5. Then, the flow matching process is safe according to Definition 4, ensuring that the generated trajectory $\mathcal{T}_{1}$ satisfies the safety condition in Definition 3.
Although valid flow matching barrier functions ensure safety for both the trajectory and the flow matching process, the selection of an appropriate regularization term $\bm{u}_{t}$ for each $t\in[0,1]$ remains a critical challenge. To minimize the influence of $\bm{u}_{t}$ on the distribution’s alignment with the target $p_{1}$, the regularization term should be kept as small as possible in magnitude. To this end, we determine $\bm{u}_{t}$ by solving the following optimization problem:
$\bm{u}_{t}=\operatorname*{arg\,min}_{\bm{u}=[{\bm{u}^{0}}^{\top},\cdots,{\bm{u}^{H}}^{\top}]^{\top}\in\mathbb{R}^{d}}\|\bm{u}\|^{2}$ (9a)
s.t. $\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{u}^{i}+\varphi(t,h(\bm{E}_{i}\mathcal{T}_{t}))h(\bm{E}_{i}\mathcal{T}_{t})\geq 0,$ $\forall i\in\{0,\cdots,H\}.$ (9b)
Given that $\|\bm{u}\|^{2}=\sum_{i=0}^{H}\|\bm{u}^{i}\|^{2}$, the optimization problem can be decoupled for each $\bm{u}^{i}$ as
$\bm{u}_{t}^{i}=\operatorname*{arg\,min}_{\bm{u}^{i}\in\mathbb{R}^{d_{s}}}\|\bm{u}^{i}\|^{2}$ (10a)
s.t. $\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{u}^{i}+\varphi(t,h(\bm{E}_{i}\mathcal{T}_{t}))h(\bm{E}_{i}\mathcal{T}_{t})\geq 0,$ (10b)
which can be consider as a quadratic programming (QP) problem due to the quadratic objective function and linear constraint w.r.t $\bm{u}^{i}$. To alleviate the derivation of the solution of this QP problem, we first introduce two auxiliary variables as follows
$\bm{b}_{t}^{i}=\left(\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\right)^{\top},$ (11)
$a_{t}^{i}={\bm{b}_{t}^{i}}^{\top}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\varphi(t,h(\bm{E}_{i}\mathcal{T}_{t}))h(\bm{E}_{i}\mathcal{T}_{t}).$ (12)
By constructing a valid FMBF to ensure $a_{t}^{i}\geq 0$ when $\bm{b}_{t}^{i}=\bm{0}_{d_{s}\times 1}$, then the solution for such a QP problem always exists. Specifically, the explicit solution is written as
$\bm{u}_{t}^{i}=\begin{cases}\bm{0}_{d_{s}\times 1},&\text{if~{}}a_{t}^{i}\geq 0\\ -{\bm{b}_{t}^{i}a_{t}^{i}}/{\|{\bm{b}_{t}^{i}}\|^{2}},&\text{otherwise}\end{cases}$ (13)
for all $i=0,\cdots,H$ and $t\in[0,1)$. Note that the function $\varphi(t,h(\bm{E}_{i}\mathcal{T}_{t}))$ is not defined at $t=1$. Therefore, the determination method for $\bm{u}_{t}$ with $t\in[0,1)$ is not implementable. Considering $\mathcal{T}_{1}$ plays the most important role as output, For $t=1$ with $\mathcal{T}_{1}=\bm{\psi}_{1}(\mathcal{T}_{0})$, a safety filter is directly designed for $\mathcal{T}_{1}$ as
[…]
which directly ensure the trajectory safety for $\mathcal{T}_{1}$.
4.2 Safe Flow Matching under Composite Constraints
Defining a safe set using a single FMBF may be inadequate for capturing intricate or multiple safety constraints. In this subsection, we introduce a composite safe defined by $N$ differentiable continuous functions $h_{j}(\cdot):\mathbb{S}\to\mathbb{R}$ with $j=1,\cdots,N$. Then a composite safe set $\mathbb{C}_{s}$ is defined as
$\mathbb{C}_{s}=\bigcap_{j=1}^{N}\mathbb{C}_{s,j},$ $\mathbb{C}_{s,j}=\{\bm{s}\in\mathbb{S}:h_{j}(\bm{s})\geq 0\},$ $\forall j=1,\cdots,N$ (15)
We extend the definition of trajectory safety, safeFM, and FMBF from the unified case to this composite framework, as detailed below.
Definition 6 (Composite Flow Matching Barrier Function).
The function $\bm{h}(\cdot)$ is called a composite flow matching barrier function (CFMBF), if it satisfies the following conditions:
  1. 1.
    For all $i=0,\cdots,H$ and $t\in[0,1)$, there exists $\bm{u}_{t}^{i}\in\mathbb{R}^{d_{s}}$ such that
    $\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{u}_{t}^{i}+\varphi_{j}(t,h_{j}(\bm{E}_{i}\mathcal{T}_{t}))h_{j}(\bm{E}_{i}\mathcal{T}_{t})\geq 0$ (16)
    hold for all $j=1,\cdots,N$, where
    $\varphi_{j}(t,h_{j}(\bm{s}))=\begin{cases}\varphi_{0,j},&\text{if~{}}h_{j}(\bm{s})\geq 0\\ \varphi_{1,j}(t),&\text{otherwise}\end{cases}$ (17)
    with $\varphi_{0,j}\in\mathbb{R}_{+}$ and a blow-up function $\varphi_{1,j}(\cdot):[0,1)\to\mathbb{R}_{+}$ satisfying $\lim_{t\to 1^{-}}\varphi_{1,j}(t)=+\infty$.
  2. 2.
    For all $i=0,\cdots,H$ and $j=1,\cdots,N$, the terminal condition holds: $h_{j}(\bm{E}_{i}\mathcal{T}_{1})\geq 0$.
Definition 7 (Trajectory Safety with Composite FMBF).
A trajectory $\mathcal{T}$ of length $H$ is safe if satisfies the condition in Definition 3 w.r.t $h_{j}(\cdot)$ for all $i=1,\cdots,N$.
Definition 8 (Safe Flow Matching with Composite FMBF).
A flow matching process is called safe w.r.t $\bm{h}(\cdot)$, if the generated trajectory $\mathcal{T}_{1}$ at $t=1$ is safe w.r.t $\bm{h}(\cdot)$ defined by Definition 7 for any initial sample $\mathcal{T}_{0}\sim p_{0}$.
Building on the safety guarantees established for a unified safe set, we demonstrate that a composite flow matching barrier function ensures the safety of both the trajectory and the flow matching process in the composite case, as formalized below.
Theorem 3.
Suppose $\bm{h}(\cdot)$ is a valid CFMBF as defined in Definition 6. Then, the flow matching process is safe according to Definition 8, and the trajectory $\mathcal{T}_{1}$ satisfies the safety conditions in Definition 7.
Same as the seriaro in the unified contracts, we incorporate a minimal regularization term $\bm{u}_{t}$ into the flow matching process while ensuring safety under composite constraints. For SafeFM under Composite Constraints, the QP problem is formulated to determine $\bm{u}_{t}$ for any $t\in[0,1)$ as
$\bm{u}_{t}=\operatorname*{arg\,min}_{\bm{u}=[{\bm{u}^{0}}^{\top},\cdots,{\bm{u}^{H}}^{\top}]^{\top}\in\mathbb{R}^{d}}\|\bm{u}\|^{2}$ (18a)
s.t. $\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{u}^{i}+\varphi_{j}(t,h_{j}(\bm{E}_{i}\mathcal{T}_{t}))h_{j}(\bm{E}_{i}\mathcal{T}_{t})\geq 0,$ (18b)
$\forall i\in\{0,\cdots,H\},\forall j\in\{1,\cdots,N\}.$
Then the problem can be decoupled into independent subproblems for each $\bm{u}_{t}^{i}$, such that
$\bm{u}_{t}^{i}=\operatorname*{arg\,min}_{\bm{u}^{i}\in\mathbb{R}^{d_{s}}}\|\bm{u}^{i}\|^{2}$ (19)
s.t. $\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{u}^{i}+\varphi_{j}(t,h_{j}(\bm{E}_{i}\mathcal{T}_{t}))h_{j}(\bm{E}_{i}\mathcal{T}_{t})\geq 0,$ $\forall j\in\{1,\cdots,N\}$
However, with multiple linear constraints, the feasibility of this decoupled QP is not guaranteed. To ensure feasibility, we introduce relaxation terms $\delta_{j}^{i}\in\mathbb{R}_{0,+}$ and reformulate decoupled QP problem as
$\bm{u}_{t}^{i}=\operatorname*{arg\,min}_{\bm{u}^{i}\in\mathbb{R}^{d_{s}},\delta_{j}^{i}\in\mathbb{R}_{0,+}}\|\bm{u}^{i}\|^{2}+\sum_{j=1}^{N}{\delta_{j}^{i}}^{2}$ (20a)
s.t. $\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{u}^{i}+\varphi_{j}(t,h_{j}(\bm{E}_{i}\mathcal{T}_{t}))h_{j}(\bm{E}_{i}\mathcal{T}_{t})+\delta_{j}^{i}\geq 0,$ (20b)
$\forall j\in\{1,\cdots,N\}.$
Remark 1.
It is straightforward to verify that choosing sufficiently large values for $\delta_{j}^{i}$ always ensures the feasibility of the QP problem. This effectively turns the problem into a feasibility problem. For instance, one feasible solution is $\bm{u}_{i}=\bm{0}_{d_{s}\times 1}$ and $\delta_{j}^{i}=\max\{0,-\frac{\partial h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})-\varphi_{j}(t,h_{j}(\bm{E}_{i}\mathcal{T}_{t}))h_{j}(\bm{E}_{i}\mathcal{T}_{t})\}$ for all $j=1,\cdots,N$.
Remark 2.
Due to such relaxation for feasibility, the positivity for $h_{j}(\bm{E}_{i}\mathcal{T}_{1^{-}})$ may not be satisfied. In such cases, the importance of condition 2 in Definition 6 is shown, where the safe trajectory $\mathcal{T}_{1}$ is obtained by solving
[…]
If $\mathbb{C}_{s}$ is non-empty, then (21) is feasible. The non-empty $\mathbb{C}_{s}$ indicates the existence of $\bm{s}_{safe}\in\mathbb{C}_{s}\subseteq\mathbb{S}$. Then, the trajectory $\mathcal{T}_{safe}=[\bm{s}_{safe}^{\top},\cdots,\bm{s}_{safe}]^{\top}$ is safe as in Definition 8, which indicates the existence of at least one solution, i.e., $\mathcal{T}_{1}=\mathcal{T}_{safe}$, for (21).
5 Experimental Results
5.1 Planar Robot Navigation
We demonstrate the effectiveness of SafeFM in a planar robot navigation scenario. We generate category-specific trajectories within a two-dimensional environment, verifying the quality and safety of generated trajectories.
Task Definition.
This experiment evaluates the performance of our SafeFM model for planar robot navigation. Specifically, the model generates two-dimensional trajectories by numerically integrating a learned deterministic vector field from random initial points sampled from a simple prior. Trajectories are categorized into two distinct types determined, determined by navigation directions from the lower-left to upper-right and from the upper-left to lower-right, respectively. The objective is twofold:
  • Verify whether the generated trajectories start and terminate within predefined regions.
  • Assess the effectiveness of safeFM in regulating trajectories to avoid obstacles.
Data and Model.
The FM model was trained on a synthetic dataset comprising two classes of planar trajectories, each with $10,000$ samples. The FM model was trained on a synthetic dataset consisting of two classes of planar trajectories, each with 10,000 samples. Every trajectory is represented as a 100-point Bézier curve, with its start and end points sampled uniformly from circles of radius 0.2. For class 0, the start and end circles are centered at (-1,-1) and (1,1), respectively; for class 1, they are centered at (-1,1) and (1,-1). During training, the model learned to map random noise to smooth, goal-directed trajectories following specific directional distributions. At inference, the model parameters remain fixed, and trajectories are generated by numerically solving an ODE initialized from random points. The CBF module subsequently corrects trajectories to enforce safety constraints.
Evaluation Metrics.
The following metrics are used to comprehensively evaluate trajectory quality and safety:
  • Success Rate: Proportion of generated trajectories with start and end points within predefined goal regions.
  • Safe Trajectory Ratio: Proportion of trajectory waypoints satisfying safety constraints, providing further insight into the effectiveness of the CBF projection.
  • Inference Efficiency: Average computational time required to generate trajectories, indicating the model’s suitability for real-time applications.
Results and Analysis.
To illustrate the effectiveness of SafeFM in generating smooth, safe, and goal-directed trajectories for planar robot navigation, we qualitatively analyze the trajectory generation process based on benchmark experiments. We conducted extensive evaluations using $1,000$ trajectory samples per class to assess both the accuracy and safety of the generated trajectories. Starting from random initial noise patterns, SafeFM model consistently and smoothly converges to trajectories that accurately reflect the learned class-specific characteristics. These generated trajectories clearly capture the directional patterns and spatial distributions inherent in the original training dataset, demonstrating the model’s ability to learn and generalize critical trajectory features for each navigation scenario.
To further confirm endpoint accuracy, Figure 1 visualizes representative trajectories from each class, along with predefined circular regions (radius = 0.2) marking the desired start and end zones. All displayed trajectories consistently initiate and terminate within these designated goal regions, demonstrating that the FM model reliably adheres to endpoint constraints.
Figure 1: Representative trajectories with goal regions (radius = 0.2). Trajectories consistently start and end within the designated zones.
Figure 2 highlights the critical role of the CBF in trajectory safety correction.
Figure 2: Trajectory safety improvement via CBF projection. Obstacle region (radius = 0.25) is centered at the origin. Trajectories after projection avoid unsafe regions.
As shown in Table 1, the baseline FM model achieves reasonably high accuracy at both start and end positions, with endpoint success rates exceeding 86% for both classes. However, FM model generated trajectories also exhibit a notable percentage of obstacle violations 12.40% for Class 0 and 12.73% for Class 1 indicating safety concerns in cluttered environments. The SafeFM model eliminates all obstacle violations entirely across both classes. In addition, SafeFM improves both start and endpoint accuracies to 100.00%, demonstrating its capability to simultaneously enforce trajectory safety and enhance positional precision. The detailed results are summarized in Table 1.
Table 1: Trajectory Accuracy and Safety Evaluation
Method Class Start Accuracy (%) End Accuracy (%) Obstacle Violation (%)
FM 0 88.60 86.40 12.40
1 86.80 90.10 12.73
SafeFM 0 100.00 100.00 0.00
1 100.00 100.00 0.00
Overall, these quantitative results clearly demonstrate the effectiveness of SafeFM in generating accurate and collision-free trajectories, highlighting its potential for real-world deployment in planar robotic navigation scenarios.
6 Conclusion
In this paper, we tackled the critical challenge of enabling robots to generate safe and efficient trajectories in dynamic and unseen environments, a fundamental requirement for autonomous systems in real-world applications. Existing generative methods, such as diffusion models, while capable of modeling complex trajectory distributions, often incur high computational costs and lack explicit safety guarantees, limiting their practicality in safety-critical scenarios. To address these shortcomings, we proposed SafeFM, a novel motion planning framework that integrates flow matching with CBFs to ensure trajectory safety while maintaining computational efficiency. SafeFM leverages the deterministic and efficient nature of flow matching to transform samples from a simple prior distribution into goal-directed trajectories, embedding safety constraints directly into the learned flow field via CBFs. This approach guarantees that generated trajectories remain within predefined safe regions throughout the planning horizon, even in the presence of unseen obstacles or dynamic constraints. Our experimental evaluations, spanning planar robot navigation, demonstrated SafeFM’s superior performance over conventional flow model planners. Specifically, SafeFM achieved higher success rates, eliminated safety violations entirely after CBF correction, and offered significantly faster inference times, making it highly suitable for real-time robotic applications.
Despite its strengths, SafeFM has limitations that warrant consideration. The current formulation relies on predefined safety constraints, which may be challenging to specify in highly complex or unpredictable settings. Additionally, while effective in the evaluated scenarios, further validation is needed to confirm its scalability to more intricate tasks or real-world conditions with greater variability. Looking ahead, future research could focus on dynamically learning or adapting safety constraints to improve flexibility, integrating SafeFM with complementary frameworks like reinforcement learning for enhanced adaptability, or extending its application to a wider range of robotic domains.